OpenStudy (anonymous):

2. The theoretical probability of undesirable side effects resulting from taking Grebex is 2 in 19. If 570 people take Grebex to lower their blood pressure, how many will encounter undesirable side effects?

OpenStudy (anonymous):

\[2/19 = .105\] \[.105 * 570 = 7.35\] So you can estimate 8 people will likely suffer undesirable side effects.

OpenStudy (anonymous):

2/19=0.105 then 0.105*570 =60 60 people